Two point particles, one with charge +8.00 nC and the other with charge -2.00 nC, are separated by 4.00 m. The electric field in N/C midway between them is:


1
Expert's answer
2021-07-01T04:56:22-0400

The electric field at the middle of a certain distance (d=4 m) can be found as:


"\overrightarrow{E_c}=\overrightarrow{E_{1c}}+\overrightarrow{E_{2c}}=k\dfrac{q_1}{{r_{1c}}^2}\,\widehat{r_{1c}}+k\dfrac{q_2}{{r_{2c}}^2}\,\widehat{r_{2c}}"


Following this we have to set the coordinates for the system so that r1=(0,0)m and r2=(4,0)m are the position vectors for the charges q1= +8 nC and q2= -2 nC and then we find everything related to the center of the distance rc=(d/2,0)=(2,0)m:


"\overrightarrow{r_{1c}}=\overrightarrow{r_c}-\overrightarrow{r_1}=(2,0)m-(0,0)m=(2,0)m
\\ \overrightarrow{r_{2c}}=\overrightarrow{r_c}-\overrightarrow{r_2}=(2,0)m-(4,0)m=(-2,0)m
\\ r_{1c}=r_{2c}=d/2=2\,m
\\ \widehat{r_{1c}}=\cfrac{\overrightarrow{r_{1c}}}{r_{1c}}=\frac{(2,0)}{2}=(1,0)
\\ \widehat{r_{2c}}=\cfrac{\overrightarrow{r_{2c}}}{r_{2c}}=\frac{(-2,0)}{2}=(-1,0)"


With this information we substitute to find "\overrightarrow{E_c}":


"\overrightarrow{E_c}=k\dfrac{q_1}{{r_{1c}}^2}\,\widehat{r_{1c}}+k\dfrac{q_2}{{r_{2c}}^2}\,\widehat{r_{2c}}"


"\overrightarrow{E_c}=(8.988\times10^9\frac{Nm^2}{C^2})\frac{(8\times10^{-9}C)}{(2\,m)^2}(1,0)+(8.988\times10^9\frac{Nm^2}{C^2})\frac{(-2\times10^{-9}C)}{(2\,m)^2}(-1,0)"


"\overrightarrow{E_c}=(17.976)(1,0)\frac{N}{C}+(-4.494)(-1,0) \frac{N}{C}=(17.976,0)\frac{N}{C}+(4.494,0)\frac{N}{C}"


"\to \overrightarrow{E_c}=(22.47,0)\frac{N}{C} \implies {E_c}=22.47 \,\frac{N}{C}"


In conclusion, the electric field midway between the charges is Ec=22.47 N/C.


Reference:

  • Young, H. D., & Freedman, R. A. (2015). University Physics with Modern Physics and Mastering Physics (p. 1632). Academic Imports Sweden AB.